# MusicInfuser: Making Video Diffusion Listen and Dance

### v1

@@ Abstract
We introduce MusicInfuser, an approach for generating high-quality dance videos that are synchronized to a specified music track. Rather than attempting to design and train a new multimodal audio-video model, we show how existing video diffusion models can be adapted to align with musical inputs by introducing lightweight music-video cross-attention and a low-rank adapter. Unlike prior work requiring motion capture data, our approach fine-tunes only on dance videos. MusicInfuser achieves high-quality music-driven video generation while preserving the flexibility and generative capabilities of the underlying models. We introduce an evaluation framework using Video-LLMs to assess multiple dimensions of dance generation quality. The project page and code are available at https://susunghong.github.io/MusicInfuser.
Figure 1: MusicInfuser adapts video diffusion models to music, making them listen and dance according to the music. This adaptation is done in a prior-preserving manner, enabling it to also accept style through the prompt while aligning the movement to the music. Please refer to the project page videos, as the movement appears slower due to the frame sampling rate.
@@ 1 Introduction
Today’s leading AI video generation tools (e.g., Sora, Gen, Veo) produce silent videos. While it’s possible to add music after the fact, it is difficult to generate motion that’s properly synchronized with a specified music track. Some research has begun to explore joint audio-video generation, e.g., [39], but this area is in its infancy and requires training larger and more complex joint audio-video models.
@@
In this paper, we introduce an approach to adapt pre-trained text-to-video models to condition on music tracks. Our method, called MusicInfuser, generates output videos that are synchronized with the input music, with styles and appearances controllable via text prompts. We focus on the specific application of dance, i.e., generating realistic dancing figures that adjust and synchronize to the music.
@@
The automatic generation of dance movements from music presents significant challenges due to the need to simultaneously consider multiple aspects such as style, beat, and the inherently multimodal nature of dance, where multiple valid movement sequences can follow from any given pose [30]. Principles of choreography [48] have inspired computational approaches to dance generation, leading researchers to explore methods ranging from graph-based approaches [28, 12] to modern deep neural networks [50, 51, 45] to synthesize dance movements. However, traditional methods for dance synthesis have relied on motion capture data [2], which is resource-intensive, or reconstructed motions [32], which often have floating and jitter issues [2].
We take a different approach by aligning music with videos generated by pre-trained Text-to-Video (T2V) models [44]. MusicInfuser does not require motion capture or motion reconstruction, but instead simply uses dance videos for training. Concretely, we propose adapter networks consisting of music-video cross-attention and a low-rank adapter, along with a novel training methodology and layer selection strategy for the cross-attention to maintain a balanced representation between multiple modalities with significant modality gaps.
Besides endowing the video diffusion model with the ability to listen to music, MusicInfuser preserves the rich knowledge in the text modality, enabling various forms of control and providing a flexible interface for the generation process. This means users can still leverage textual prompts to guide dance style, setting, and other aesthetic elements while maintaining synchronization with music (Fig. 1). Moreover, our framework can even generalize to make group choreography (Fig. 2) and longer dance videos with unseen music tracks (Fig. 3). To systematically evaluate our results, we developed an automatic evaluation framework using Video-LLMs [31] capable of processing video, audio, and language information simultaneously. This comprehensive approach allows us to assess multiple dimensions of dance quality, video quality, and prompt alignment within a single framework.
@@
Our experiments demonstrate that MusicInfuser effectively bridges the gap between music and videos without requiring specialized motion data. By leveraging existing video diffusion models through targeted adaptation, we achieve high-quality and novel dance movements that respond naturally to musical rhythms and patterns, offering versatility in response to prompts and suggesting another direction for music-driven choreography synthesis.
@@ 2 Related Work
Early music-to-dance generation research developed frameworks mapping music primitives to dance elements, using Hidden Markov Models [35]. Graph-based methods built movement transition graphs synchronized to musical beats [28], using quality rating functions and constraint-based dynamic programming [15]. Later, researchers incorporated Gaussian processes [16], Conditional Restricted Boltzmann Machines, Recurrent Neural Networks [1], and Convolutional Neural Networks [50, 51]. More recent approaches incorporate transformers [50, 32]. For instance, a Full-Attention Cross-Modal Transformer [32] predicts dance sequences based on seed motion and audio information. Traditional approaches typically generated movements synchronized with beats but lacking contextual meaning or showing excessive repetition [3], while showing limited choreographic diversity [4] and struggling with generalization. Recent dance generation advances have shifted toward diffusion-based approaches to address limitations of earlier methods [36, 2, 45, 37, 29]. Unlike these approaches that focus on motion skeleton synthesis from music, MusicInfuser directly synthesizes dance videos and choreography by uniquely adapting pre-trained text-to-video diffusion models to incorporate musical inputs while preserving their inherent knowledge of diverse dance styles and general human movements.
@@
Dance generation systems have evolved to incorporate multiple input modalities for richer choreographic control [8, 34, 18], with text emerging as a powerful interface for its zero-shot capability and communicating choreographic ideas [34]. Transformer-based approaches using Vector Quantized-Variational Autoencoders create discrete motion tokens processable alongside text [41], while systems now process both text and music inputs simultaneously [18]. Our MusicInfuser framework combines the flexibility of text-based interfaces with precise audio synchronization, allowing users to control stylistic and aesthetic elements of generated dance videos through prompts while ensuring movements remain aligned with musical features.
@@
Figure 2: Due to the conservation of knowledge in video and text modalities, our model generalizes to generate group dance videos by modulating the prompt. To show this, the prompt is set to [DANCERS] dancing in a studio with a white backdrop, captured from a front view,” where [DANCERS] denotes the description for each number of dancers.
Figure 3: We generate longer dance videos that are 2 times longer than the videos used for training. For each row, we use synthetic in-the-wild music tracks with a keyword “K-pop,” a type of music not existing in AIST [46], and use a prompt “a professional dancer dancing K-pop ….” This shows our method is highly generalizable, even extending to longer videos with an unheard cateory of the music. The beat and style alignment can be more clearly observed in the videos on the project page.
@@ 3 Preliminaries
$\frac{d\mathbf{x}}{d\sigma}=$ $-\gamma_{\text{cfg}}\left[\frac{D_{\theta}(\mathbf{x}|\mathbf{c};\sigma)-\mathbf{x}}{\sigma}\right]$ (3)
$+(\gamma_{\text{cfg}}-1)\left[\frac{D_{\theta}(\mathbf{x};\sigma)-\mathbf{x}}{\sigma}\right].$ (4)
@@ 4 MusicInfuser
These models, trained on diverse video datasets, have already internalized a wide range of human motion patterns including dance movements. The choreographic knowledge encoded in these models represents a valuable foundation that can be leveraged for music-driven video generation. Therefore, our goal is to adapt the models to construct a final denoiser, $D_{\theta}(\mathbf{x}|\mathbf{c},\mathbf{a};\sigma)$, that is also conditioned on a music condition $\mathbf{a}$. Rather than teaching these models to dance from scratch, our approach focuses on effectively aligning this pre-existing knowledge with audio input. In this section, we address methods that preserve the rich knowledge embedded in text-to-video models while establishing correlations between musical features and dance movements. The illustration of the overall architecture is displayed in Fig. 4.
4.1 Zero-Initialized cross-attention
Attention mechanisms [47] provide an effective mechanism for conditioning diffusion models [38, 40]. The cross and self attention are known to highly contribute to the structure of the generated images or videos [13, 23, 26, 40, 10, 19]. In our framework, we propose a Zero-Initialized cross-attention (ZICA) adapter to condition the musical information of the audio, while preserving the original capabilities of the base model.
Specifically, the audio signal is first processed through a specialized audio encoder [5] that extracts relevant temporal and spectral features. These encoded audio representations then pass through a learnable projector that maps them into the same embedding space as the video tokens. Then, the projected audio tokens interact with video tokens through cross-attention layers, allowing the model to establish correlations between audio patterns and corresponding visual choreography. Formally, if $\mathbf{A}$ denotes the audio embeddings and $\mathbf{V}$ denotes video tokens, the cross-attention operation can be expressed as:
$\text{Attention}(\mathbf{V},\mathbf{A})=\text{softmax}\left(\frac{\mathbf{W}_{Q}\mathbf{V}\cdot(\mathbf{W}_{K}\mathbf{A})^{T}}{\sqrt{d}}\right)\mathbf{W}_{V}\mathbf{A}.$ (5)
The complete cross-attention mechanism includes an output transformation matrix $\mathbf{W}_{O}$ that projects the attention output back to the original feature space, with the full operation being:
$\mathbf{Z}=\mathbf{V}+\mathbf{W}_{O}\text{Attention}(\mathbf{V},\mathbf{A}).$ (6)
By initializing $\mathbf{W}_{O}$ to zero during the early stages of training, we effectively disable the influence of audio conditioning on the video features. This initialization strategy ensures that $\mathbf{Z}=\mathbf{V}$, meaning the output remains identical to the input video features regardless of audio content. As training progresses, the parameters of $\mathbf{W}_{O}$ gradually move away from zero, allowing the model to incrementally incorporate audio-conditioned information. This controlled integration preserves the rich choreographic knowledge of the base model while allowing for the gradual emergence of audio-visual correlations without disrupting the foundational video generation capabilities.
Figure 4: Overall architecture of MusicInfuser. Our framework adapts a pre-trained diffusion model with audio embedding using ZICA blocks (Sec. 4.1) and HR-LoRA blocks (Sec. 4.2). The placement of ZICA blocks is selected based on layer adaptability (Sec. 4.6).
4.2 Low-Rank Adaptation with Higher Rank (HR-LoRA)
In addition to the cross-attention, we adapt the attention weights for diffusion transformer blocks. The adapter serves two key purposes: (1) to effectively integrate audio features into the text-video processing pipeline, and (2) to shift the domain toward our target application, synthesizing clear choreography. The ranks conventionally adopted for Low-Rank Adaptation (LoRA) [24] for visual models (typically 8 or 16) are optimized for image models and often lead to failure in capturing the full complexity of spatiotemporal information when directly applied to video models.
For adapting video tokens, a higher rank is required compared to image tokens, since video tokens contain temporal information. To effectively model motion adaptation separately from spatial adaptation, the optimal rank for the linear map should be increased compared to what’s needed for static images. For instance, adapting the full dimension of homography requires increasing the optimal rank by at least 8 (the degree of freedom for homography), while adapting general videos or complex human motion necessitates even higher ranks.
@@ 4.3 Beta-Uniform Scheduling
Current diffusion models, including those using LoRA fine-tuning, typically employ uniform scheduling for noise sampling during training. However, for adapter training for dance video generation, our aim is to preserve the pre-trained models’ denoising capability of major components like coarse human motion at early stages and gradually learn the components over the course of the training process. To this end, we propose a Beta-Uniform scheduling strategy that makes the training noise distribution $\Sigma_{\text{train}}$ evolve from a Beta distribution to a uniform distribution.
@@
This causes a smooth transition from focusing on high-frequency components at lower noise levels to broadly considering all frequencies. By first influencing the task-specific fine components of the dance and then the fundamental structure of dance movements, our approach preserves the rich knowledge of human motion and produces more coherent dance sequences. See the supplementary material for the details.
Figure 5: Comparison of audio-driven generation with MM-Diffusion [39]. Our method produces fewer artifacts (shown in the first and third rows), while generating more realistic dance videos with more natural movements (first row) and more dynamic motion (second and third rows). Note that we use the same music track for each row, and the spectrogram is stretched for MM-Diffusion since we generate longer videos. For our method, we use the fixed caption “a professional dancer dancing …” across all music tracks.

### v2

@@ Abstract
We introduce MusicInfuser, an approach that aligns pre-trained text-to-video diffusion models to generate high-quality dance videos synchronized with specified music tracks. Rather than training a multimodal audio-video or audio-motion model from scratch, our method demonstrates how existing video diffusion models can be efficiently adapted to align with musical inputs. We propose a novel layer-wise adaptability criterion based on a guidance-inspired constructive influence function to select adaptable layers, significantly reducing training costs while preserving rich prior knowledge, even with limited, specialized datasets. Experiments show that MusicInfuser effectively bridges the gap between music and video, generating novel and diverse dance movements that respond dynamically to music. Furthermore, our framework generalizes well to unseen music tracks, longer video sequences, and unconventional subjects, outperforming baseline models in consistency and synchronization. All of this is achieved without requiring motion data, with training completed on a single GPU within a day.
Figure 1: MusicInfuser adapts video diffusion models to music, making them listen and dance according to the music. This adaptation is done in a prior-preserving manner, enabling it to also accept style through the prompt while aligning the movement to the music.
Figure 2: Motivational example. Skeletal motion generation [44] produces simplified movements lacking nuances such as backbone curvature, axial rotation, hand articulation, hair dynamics, and clothing motion, resulting in a more limited range of dance compared to video-based dance generation approaches (ours).
@@ 1 Introduction
Today’s leading AI video generation tools (e.g., Sora, Gen, Veo) often produce silent videos. While it is possible to add music after the fact, it is difficult to generate motion that is properly synchronized with a specified music track. Alternatively, some research has begun to explore audio-video generation [38]. However, focusing on the specific application of dance, dance videos well-aligned with their music are far rarer than finding general, unconstrained videos, resulting in sub-optimal quality when training audio-video generative models from scratch.
@@
In this paper, we introduce an approach to align pre-trained text-to-video models that have useful ingredients for dance. Our method, called MusicInfuser, generates output videos that are synchronized with the input music, with various components such as styles and appearances controllable via text prompts. We focus on synthesizing dance videos, i.e., generating realistic dancing figures that adjust and synchronize to the music, which poses several difficulties that require extensive knowledge about human motion and physics, music, and choreography.
@@
Automatic dance generation must consider style, beat, and the inherently multimodal nature of dance, where multiple valid sequences can follow a given pose [30]. Computational approaches have drawn on choreographic principles [47] and techniques ranging from graph-based methods [28, 12] to deep neural networks [50, 51, 44]. However, traditional methods rely on motion capture [3] or reconstructed motions [32], which are costly or prone to floating/jitter artifacts. In addition, skeletal representations are underparameterized for dance, lacking nuances such as backbone curvature, axial rotation, hand articulation, hair dynamics, and clothing motion (Fig. 2).
MusicInfuser bypasses these limitations by adapting pre-trained text-to-video models [43] with zero-initialized music-video modules injected into DiT blocks. This approach does not require motion capture or reconstruction, relying instead on existing dance videos for alignment. To address the scarcity of high-quality music-aligned dance datasets and reduce fine-tuning costs, we introduce a layer-wise adaptability criterion using a guidance-based constructive influence function. This preserves pre-trained knowledge while establishing correlations between music and movement, allowing training on a GPU within a day.
MusicInfuser retains text-based control, enabling users to specify dance style, setting, and other aesthetics (Fig. 1) as well as the number of dancers (Fig. 4) while maintaining music synchronization. Our method generalizes to longer videos with unseen music and even to unseen subjects such as animals (Figs. 36). For evaluation, we introduce an automatic framework based on Video-LLMs [14, 46] that jointly assesses video, audio, and language alignment, correlating well with human judgment.
@@
Our experiments show that MusicInfuser successfully closes the gap between music and dance without intermediate motion data. By leveraging pre-trained video diffusion models through targeted adaptation, it produces high-quality, novel dance movements that respond naturally to musical rhythms, offering flexible dance video generation.
Figure 3: Using prompts such as “a {marmot, rabbit, dog (top to bottom rows)} dancing …,” our method generalizes to unseen dancing subjects.
@@ 2 Related Work
Figure 4: We can generate group dance videos aligned with music, based on the text.
@@
Early approaches mapped music primitives to dance elements using Hidden Markov Models [35] and graph-based methods with movement transition graphs [28]. Later research integrated Gaussian processes [15], various neural networks [2, 50, 51], and transformers [50, 32]. Traditional methods often produced beat-synchronized movements lacking contextual meaning or showing excessive repetition [4], with limited choreographic diversity [5]. Recent advances have shifted toward diffusion-based approaches [36, 3, 44, 37, 29]. Unlike these skeleton-based methods, our framework directly synthesizes dance videos by adapting pre-trained text-to-video diffusion models to musical inputs. Without an intermediate representation, our method avoids rigid body parameterization, requires no motion capture or pose reconstruction, and eliminates post-processing to generate dance videos.
@@
Dance generation systems have evolved to incorporate multiple input modalities for richer choreographic control [9, 34, 17], with text emerging as a powerful interface for its zero-shot capability and communicating choreographic ideas [34]. Transformer-based approaches using Vector Quantized-Variational Autoencoders create discrete motion tokens processable alongside text [40], while systems now process both text and music inputs simultaneously [17]. The MusicInfuser framework combines the flexibility of text-based interfaces with precise audio synchronization, allowing users to control stylistic and aesthetic elements of generated dance videos through prompts while ensuring movements remain aligned with musical features.
@@
Figure 5: Zero-initialized cross-attention (ZICA) block. The output projection is initialized with a zero matrix, making the cross-attention block act as an identity function at the beginning. (b–d) illustrate several baseline layer selection strategies when the number of ZICA blocks is fewer than that of DiT blocks. (b) Attaching cross-attention blocks evenly across DiT blocks. (c) Attaching the blocks evenly to the earliest layers. (d) Attaching the blocks based on pre-computed layer adaptabilities (Sec. 4.1). Table 1 shows the results.
@@ 3 Preliminaries
$\frac{d\mathbf{x}}{d\sigma}=-\gamma_{\text{cfg}}\left[\frac{D_{\theta}(\mathbf{x}|\mathbf{c};\sigma)-\mathbf{x}}{\sigma}\right]+(\gamma_{\text{cfg}}-1)\left[\frac{D_{\theta}(\mathbf{x};\sigma)-\mathbf{x}}{\sigma}\right].$ (3)
@@ 4 MusicInfuser
In contrast to previous multimodal dance generation methods, video diffusion models trained on massive and diverse video datasets have already internalized sophisticated representations of human motion, including intricate and expressive movements. They have implicitly learned choreographic patterns, style variations, and the general physics of a human body during their extensive training, providing a valuable foundation that can be leveraged for music-driven video generation.
Considering this, our goal is to align the models to musical input $\mathbf{a}$ with adaptation parameters $\phi$ to construct a final denoiser, $D_{\theta,\phi}(\mathbf{x}|\mathbf{c},\mathbf{a};\sigma)$, while preserving the pre-trained model’s knowledge about dance. For the rest of this paper, we call the probability distribution characterized by the pre-trained text-to-video denoiser $D_{\theta}(\mathbf{x}|\mathbf{c};\sigma)$ the prior, since it denotes a learned prior video distribution not conditioned on audio. Accordingly, our new continual optimization objective is as follows:
$\mathcal{L}=\mathbb{E}_{\mathbf{(y,c,a)}\sim p_{\textrm{mm}},\sigma\sim\Sigma_{\text{train}},\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})}\|D_{\theta,\phi}(\mathbf{y}+\mathbf{n}|\mathbf{c},\mathbf{a};\sigma)-\mathbf{y}\|^{2}_{2},$
where $p_{\textrm{mm}}$ is a joint data distribution of video, caption, and audio.
Unfortunately, we recognize that specialized dancing datasets are notably scarcer than general video datasets used for pre-training and thus inevitably contain biases that compromise the prior model’s generalization and denoising capabilities. Moreover, significant resources are required for fine-tuning text-to-video diffusion models. We address this challenge through a carefully balanced adaptation mechanism that preserves the rich prior while establishing robust correlations between musical features and dance movements with a significantly lower cost.
Figure 6: Generalization capabilities in terms of music length and type. MusicInfuser can generate multiple times longer dance videos that are multiple times longer than the videos used for training. For each row, we use synthetic in-the-wild music tracks with a keyword “K-pop,” a type of music not existing in AIST [45], and use a prompt “a professional female dancer dancing K-pop ….” This shows our method is highly generalizable, even extending to longer videos with an unheard cateory of the music. The beat and style alignment can be more clearly observed in the supplementary video.
4.1 Measuring Layer Adaptability
Cross-attention is effective for conditioning on auxiliary modalities [13, 39, 18]. However, applying cross-attention mechanisms to all layers of a model poses challenges: 1) it incurs substantial computational costs, and 2) it can degrade the denoising capabilities of pre-trained diffusion models (see “All Layers” in Table 1), especially in low-data regimes such as professional dancing. Consequently, identifying an optimal subset of layers for cross-attention adaptation is important to preserve both the denoising effectiveness and the generalization capability of the pre-trained model.
Unfortunately, finding this optimal combination through exhaustive search by fine-tuning every possible configuration is infeasible. For example, for a pre-trained model with 48 layers, the number of possible combinations for adapting only one-third of the layers with cross-attention is $\binom{48}{16}>2\times 10^{12}$. Two intuitive layer selection approaches, shown in (b) and (c) of Fig. 5, reduce computational costs. However, both methods fail to account for the behavior of the layers and compromise the model’s capabilities, specifically degrading video outputs (Table 1).
To address this, we propose a principled, constructive metric for layer selection based on adaptability. Instead of measuring importance by performance degradation when a layer is removed, we measure each layer’s positive influence by using it as guidance during inference. This way, we can use existing evaluation metrics for videos [25] to measure and precompute the influence of each layer without the risk of out-of-distribution issues. Specifically, this criterion quantifies each layer’s influence by performing guided sampling while leveraging the pre-trained model without the layer to provide guidance [26, 22, 1]. The layer skip guidance can be formulated as the derivative of the implicit energy function [26, 23]:
[…]
Figure 7: Speed control. The audio input is slowed down (top row) or sped up (bottom row) by factors of 0.75 and 1.25, respectively. This shows that speeding up audio generally results in sped-up movement. Note also the change in dynamics, as speeding up the audio increases the musical tone. More examples of audio speeding up and slowing down are included in the supplementary video.
Ours 8.95 9.54 10.00 7.36 5.25 8.22 7.08 7.01 9.96 8.02 8.14
All Layers 8.37 9.02 9.55 7.02 5.35 7.86 6.00 7.17 9.95 7.71 7.80
Evenly Distributed Layers 8.15 9.01 9.95 6.59 5.08 7.76 5.93 6.61 9.66 7.40 7.62
First Layers 8.67 9.44 9.90 7.20 5.57 8.16 6.36 6.88 9.89 7.71 7.99
Middle Layers 9.05 9.39 9.05 6.67 5.56 7.94 5.84 6.80 9.78 7.47 7.77
Last Layers 8.60 9.34 9.80 6.91 5.45 8.02 6.03 6.70 9.81 7.51 7.83
Feature Addition 8.60 9.37 9.90 6.92 5.41 8.04 6.29 6.90 9.94 7.71 7.92
No Beta-Uniform Schedule 8.93 9.42 9.40 6.67 5.61 8.01 6.20 6.99 9.93 7.71 7.89
No In-the-Wild Data 8.80 9.52 9.75 7.17 5.39 8.13 6.46 6.56 9.96 7.66 7.95
No Cross Attention Zero Init. 8.64 9.25 9.95 7.05 5.28 8.03 6.56 6.91 9.98 7.82 7.95
Table 1: Evaluation with Qwen3-Omni [46] shows various baseline cross-attention adaptation strategies. Feature Addition refers to directly adding audio features, inspired by image conditioning in ControlNet [49]. Results using VideoLLaMA 2 [14] are provided in the supplementary material, which show a similar trend in layer selection.
@@ 4.2 Beta-Uniform Scheduling
Diffusion models, including those using LoRA fine-tuning, typically employ a uniform distribution for noise sampling throughout training. For adapter training, we aim to preserve the denoising capability of the pre-trained model by initially focusing on low-noise levels and gradually learning the more substantial components over the course of training. To achieve this, we propose a Beta-Uniform scheduling strategy that evolves the training noise distribution $\Sigma_{\text{train}}$ from a Beta distribution concentrated on low noise levels to a uniform distribution.
@@
This causes a smooth transition from focusing on high-frequency components at lower noise levels to equally considering all frequencies. By first influencing the task-specific fine components of the dance and then the fundamental structure of dance movements, our approach preserves the pre-trained knowledge of general physics of human motion and produces more coherent dance sequences.
4.3 Zero-Initialized Adaptation Modules
To extend pre-trained diffusion transformers to new modalities while maintaining stable training, we introduce zero-initialized adaptation modules that start with zero parameters and gradually learn to influence the model. Specifically, we employ Zero-Initialized Cross-Attention (ZICA) for multimodal conditioning and Low-Rank Adaptors (LoRA) for domain and motion adaptation with the new modality.
Random initialization of cross-attention modules can bias predictions and destabilize continual training. ZICA addresses this by initializing the output projection to zero, so that cross-attention initially behaves as an identity mapping and gradually incorporates information from the conditioning modality. Let $\mathbf{A}\in\mathbb{R}^{N^{A}\times d}$ and $\mathbf{V}\in\mathbb{R}^{N^{V}\times d}$ denote projected audio and video tokens. The cross-attention with output projection is
$\mathbf{Z}=\mathbf{V}+\mathbf{W}_{O}\,\text{softmax}\left(\frac{\mathbf{V}\mathbf{W}_{Q}(\mathbf{A}\mathbf{W}_{K})^{\top}}{\sqrt{d}}\right)\mathbf{A}\mathbf{W}_{V},$ (6)
where $\mathbf{W}_{O}$ is initialized to zero. The module initially acts as an identity mapping, and as $\mathbf{W}_{O}$ moves away from zero during training, it gradually integrates audio features.
Similarly, we employ LoRA [24] for two purposes: (1) adapting the model to the new domain, and (2) enabling the network to process the additional audio information introduced by ZICA. LoRA decomposes weight updates into low-rank matrices (i.e., $\Delta W=BA$) with $B$ initialized to zero, so the effective update begins from a neutral state and gradually learns task-specific modifications. While conventional LoRA ranks (e.g., 8–16) are often sufficient for image models, video transformers, which must capture temporal dependencies and cross-modal interactions, benefit from higher-rank configurations, since temporal and modality adaptation typically require more expressive updates. In particular, modeling complex human motion or scene transformations over time may demand greater capacity; we therefore adopt a higher rank.